OpenStudy (anonymous):

Use a system of equations to solve the following problem: Charles High School has a total of 94 boys and girls who play sports. If the number of boys is 22 more than twice the number of girls, how many boys play sports at this high school? 70 48 24 80

OpenStudy (anonymous):

b+g = 94 b-22 = 2g ---> b-2g = 22

OpenStudy (anonymous):

so solve the system of equations b+g = 94 b-2g = 22

OpenStudy (anonymous):

that's my problem i don't know how to do that

OpenStudy (anonymous):

okay it is just like the last problems, we need to get rid of g so that we can solve for b

OpenStudy (anonymous):

we can do that by multiplying the top equation by 2, so that the coefficient of g on both equations are opposite 2b +2g = 188 b-2g = 22 add the equations together 2b+2g = 188 + b - 2g = 22 _______________ 3b = 210 b = 70

OpenStudy (anonymous):

ohhhhhhh

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

you're welcome